OpenStudy (anonymous):

a baker misread the directions and used 4 3/4 cups of flour in a recipe. it was 1 4/9 to much flour. how much flour should he have used?

OpenStudy (anonymous):

4 + 3/4 = 19/4 1 +4/9 = 13/9 19/4-13/9 = 171-52/36 = 119/36 = 3.305 cups